A solution of Gulcose in water leveled as 10% weight by weight . The density of solutions is 1.20g /ml. Calculate the molality

Grade:12

1 Answers

PASUPULETI GURU MAHESH
7 years ago
Given solute is glucose(c6h1206)   (m.w=180 g/mol)    and solvent is water.
weight of glucose is 10g in 100g 0f water. and also  1g of water==1ml
then 10g of glucose is 10g in 100ml 0f water.
then density=mass /volume
mass= density*volueme=1.20*1000=1200
mass of the solution=1200g
for 100 ml 10g dissolved
for 1000ml 100g dissolved
then mass of solvent= mass of solution-mass of the solute
                                  =1200-100
                                  =1100g
now molality= (weight of solute in 1000ml *1000)/(molecular weight of solute*mass of solvent)
                      =(100*1000)/(180*1100)
                      =0.5050 m
   thank you